Tom is 5 times as old as Tod. In 6 years Toms age will be 8 years more than 3 times as old as tod. How old is tod?

Answer by stanbon(75887)

Tom is 5 times as old as Tod. In 6 years Toms age will be 8 years more than 3 times as old as tod. How old is tod?
---------
Equation:
m = 5d
(m+6) = 3(d+6)+8
-----
Substitute for "m" and solve for "d":
5d+6 = 3d + 26
2d = 20
d = 10 (Tod's age now)
=========================
cheers,
Stan H.
===============